# Position Paper: An Integrated Perspective on Data, Metrics, and Methodology for Deep Time-Series Forecasting

### v3

@@ Abstract
Deep time-series forecasting plays an integral role in numerous practical applications. However, existing research fall short by focusing narrowly on either neural architecture designs for long-term point forecasts or probabilistic models for short-term scenarios. By proposing a comprehensive framework, facilitated by a novel tool, ProbTS, that integrates diverse data scenarios, evaluation metrics, and methodological focuses, we aim to transcend the limitations of current forecasting practices. Rigorous experimentation uncovers pivotal insights, including the supreme importance of aligning forecasting methodologies with the unique characteristics of the data; the necessity of a broad spectrum of metrics for accurately assessing both point and distributional forecasts; and the challenges inherent in adapting existing forecasting methods to a wider range of scenarios. These findings not only challenge conventional approaches but also illuminate promising avenues for future research, suggesting a more nuanced and effective strategy for advancing the field of deep time-series forecasting. [^1] The ProbTS tool will be open-sourced.
time-series forecasting, point forecasting, probabilistic forecasting, neural networks
@@ 1 Introduction
Deep time-series forecasting serves a crucial role in a myriad of practical applications, from traffic flow forecasting (Lv et al., 2014) and renewable energy forecasting (Wang et al., 2019), to diverse forecasting demands in retail (Böse et al., 2017), finance (Hou et al., 2021), and climate (Mudelsee, 2019). Each of these domains presents unique data characteristics, such as trend movements (Taylor & Letham, 2018), periodic patterns (Smyl, 2020), complicated mixtures of global seasonality and local variations (Fan et al., 2022), as well as intricate distributional dependencies (Rasul et al., 2021b). These data characteristics collectively pose substantial challenges for accurate forecasting, and thereby necessitate the development of sophisticated modeling designs. In addition to accurate point forecasts, modern decision-making processes often require robust distributional forecasts to effectively manage uncertainty (Gneiting & Katzfuss, 2014; Hyndman & Athanasopoulos, 2018). These collective requirements emphasize the significance for methodologies that are not only versatile across various forecasting scenarios but also capable of delivering both point and distributional forecasts with high fidelity.
However, these practical needs also highlight a distinct gap in existing research for deep time-series forecasting, which tends to bifurcate into two streams: one focusing on neural architecture designs optimized for point forecasting in long-term scenarios, and the other on advancing probabilistic forecasting, primarily in short-term situations.
Particularly, one stream of research, exemplified by studies such as (Zhou et al., 2021; Wu et al., 2021; Liu et al., 2022; Zhang et al., 2022; Challu et al., 2023; Wu et al., 2023; Nie et al., 2023), primarily explores neural architecture designs tailored for time-series forecasting. These studies, however, mostly focused on point forecasts, and their architectural designs are largely driven by long-term forecasting scenarios that exhibit strong trending and seasonal patterns. Therefore, it remains unclear how these advancements in neural architecture designs can be effectively extended to distributional forecasts, and whether their specialized designs, which are closely tied to certain data patterns, maintain their effectiveness across varied data scenarios.
On the other hand, another line of research, including works like (Rasul et al., 2021b, a; Tashiro et al., 2021; Bilos et al., 2023; Kollovieh et al., 2023), emphasizes the adaption of deep generative modeling (Dinh et al., 2017; Papamakarios et al., 2017; Ho et al., 2020) to probabilistic time-series forecasting. Despite significant progress in characterizing complex data distributions without the need for a predefined closed-form prior distribution such as Gaussian (Salinas et al., 2019), these probabilistic models have primarily been developed and evaluated in short-term forecasting scenarios. Hence, it remains to be seen whether these designs can be effectively transitioned to long-term forecasting, and whether improvements in distributional forecasts also yield benefits in point forecasts, especially when compared to the tailored architectural designs mentioned earlier.
In a bid to systematically bridge this research gap, in this paper, we propose a holistic framework that integrates diverse data scenarios, comprehensive evaluation metrics, and methodological innovations. This approach enables a nuanced analysis of why certain forecasting methodologies outperform others in specific contexts and how they could be optimized to improve both point and distributional forecasting. To be more concrete, firstly, in terms of data, we encompass various data scenarios, spanning different domains and forecasting horizons. More significantly, we aim to quantify key data characteristics, elucidating why certain methodological choices converge or diverge. Secondly, concerning metrics, we adopt a comprehensive set of evaluation strategies, assessing the performance of both point and distributional forecasts. Lastly, we revisit a selection of representative studies from different research branches, affirming their advancements in specific methodological focuses and examining their performance in seldom evaluated scenarios. Through this integrated lens, we aspire to illuminate previously unaddressed questions, expose overlooked limitations of existing methods in certain situations, and reveal our novel findings that could potentially open up new avenues for future research.
To achieve these ends, we develop a tool, ProbTS, which incorporates customized data, evaluation, and model modules in alignment with the integrated perspective. Rigorous experimentation with this tool has led us to a number of intriguing insights, which are concisely presented below and explored in greater detail in Section 4.
Firstly, regarding the data, we have discovered that the distinctive characteristics of data in different forecasting scenarios primarily dictate the selection of methodology. We have identified several fundamental indicators that assist in distinguishing the distinct needs of diverse forecasting scenarios, such as data from various domains, short-term versus long-term forecasting, and point versus distributional forecasting. Notably, these indicators also facilitate our understanding of why certain methodological choices align or diverge in particular cases.
Secondly, concerning metrics, we have discerned that the use of a comprehensive suite of metrics is critical for both point and distributional forecasts. This conclusion stems from our observation that optimal performance in distributional metrics does not necessarily translate to optimal performance in point forecasts, and vice versa. Importantly, this suggests that focusing solely on optimizing one type of evaluation objective could result in sub-optimal performance according to other evaluation criteria.
Thirdly, in relation to the first research branch primarily dedicated to developing new architectures tailored for long-term point forecasting scenarios, we have recognized significant challenges and unresolved questions when adapting them to new scenarios, particularly those with distributional or short-term requirements. The integration of these architectures with advanced distribution estimation capabilities remains an open issue. Furthermore, the advantages of specific designs created for long-term scenarios diminish significantly in short-term cases. Another interesting observation is that all these methods consistently favor non-autoregressive decoding schemes over their autoregressive counterparts for multi-step forecasting, likely due to the pronounced trends in many long-term forecasting scenarios.
In terms of the second research branch, which focuses on developing advanced probabilistic forecasting capabilities, primarily in short-term scenarios, we have observed that extending existing probabilistic models to long-term scenarios presents significant challenges. Specifically, they often underperform those point forecasting methods even in producing distributional forecasts. Different from the first branch, these probabilistic methods exhibit a balanced preference for autoregressive and non-autoregressive decoding schemes. Our analyses reveal that autoregressive probabilistic methods tend to propagate more errors as the forecasting horizon increases, and non-autoregressive probabilistic methods demand substantial computational resources, both in terms of memory and time, as the forecasting horizon expands significantly. Interestingly, we also discovered that autoregressive probabilistic methods exhibit significant advantages in dealing with strong seasonality, even in long-term forecasting scenarios. This suggests that autoregressive decoding holds substantial potential for long-term forecasting, provided we can effectively mitigate the error-propagation effect amplified by pronounced trending.
Lastly, by leveraging our versatile tool that facilitates a deeper understanding of the interplay between data characteristics, forecasting performance, and methodological choices, we underscore several under-explored yet valuable directions that not only illuminate the path forward for academic research but also hold profound implications for practical applications in a multitude of sectors.
@@ 2 Related Work
As outlined in the introduction, deep time-series forecasting can be divided into two primary branches. This section provides a more thorough review of these branches, as well as a brief overview of existing forecasting toolkits. Besides, we include a refined comparison in Appendix A.1.
Neural Architecture Designs in Time-series Forecasting.
Significant research efforts have been devoted to improving neural architecture designs for time-series forecasting, with contributions ranging from extensions of multi-layer perceptrons (Oreshkin et al., 2020; Zhang et al., 2022) to customized recurrent or convolutional neural networks (Lai et al., 2018; LIU et al., 2022), and Transformer-based variants (Vaswani et al., 2017; Zhou et al., 2022). These studies typically employ non-autoregressive decoding schemes and focus on long-term forecasting scenarios, often characterized by strong trend and seasonality patterns. However, they predominantly limit themselves to point forecasts, capturing only the average future variations. While approaches like quantile regression (Wen et al., 2017; Lim et al., 2021) can mitigate this limitation, they still cannot replace the inherent capture of the data distribution.
Probabilistic Estimation in Time-series Forecasting.
In contrast, the second branch of research, known as deep probabilistic time-series forecasting, focuses on leveraging deep neural networks to capture the complex data distribution of future time series. These studies span from early approaches using pre-defined likelihood functions (Rangapuram et al., 2018; Salinas et al., 2020), to Gaussian copulas (Salinas et al., 2019; Drouin et al., 2022), and to more recent methods exploiting advanced deep generative models (Rasul et al., 2021b; Bilos et al., 2023). While providing probabilistic forecasts, they mainly evaluate short-term forecasting scenarios. Interestingly, unlike the first branch, this branch utilizes both autoregressive (Rasul et al., 2021b, a) and non-autoregressive (Tashiro et al., 2021; Bilos et al., 2023; Kollovieh et al., 2023) decoding schemes. Furthermore, this branch tends to use general neural architectures instead of customizing them. While a few methods (Tashiro et al., 2021; Li et al., 2022; Bergsma et al., 2022) incorporate customized network designs, more studies (de Bézenac et al., 2020; Rasul et al., 2021b, a; Bilos et al., 2023; Drouin et al., 2022) leverage standardized neural networks to encode time-series representations.
@@
We observe a plethora of toolkits that have been developed for time-series forecasting. These range from those primarily designed for point forecasting, such as Prophet (Taylor & Letham, 2018), sktime (Löning et al., 2019), tsai (Oguiza, 2022), and TSlib (Wu et al., 2023), to others that incorporate probabilistic forecasting, including GluonTS (Alexandrov et al., 2020), PyTorchTS (Rasul et al., 2021b), PyTorchForecasting222github.com/jdb78/pytorch-forecasting, and NeuralForecast333github.com/Nixtla/neuralforecast. In the development of ProbTS, we drew upon these existing tools, particularly PyTorchTS and GluonTS. However, ProbTS distinguishes itself by its specialized design towards the integrated perspective proposed in this paper, encompassing modularized designs across data, methodology, and evaluation components. Furthermore, we incorporate more state-of-the-art forecasting models in ProbTS, by migrating their official implementations and conducting extensive hyper-parameter tuning to ensure optimal performance.
@@ 3 The ProbTS Tool
We aggregate publicly accessible datasets used for both short-term and long-term forecasting. Initial data visualization analyses reveal that the data domains and forecasting horizons significantly influence specific data characteristics within a given forecasting horizon. For instance, many long-term forecasting scenarios exhibit clear trend and seasonality patterns within a forecasting window, while numerous short-term forecasting cases display irregular variations within a short sliding window. Consequently, we have developed quantified indicators, such as trend and seasonality strengths, along with non-Gaussianity to indicate the complexity of data distribution within a forecasting window. Detailed information about dataset statistics, visualization analyses, and quantified measures can be found in Appendix A.2A.3A.4, and A.5. The quantified measurements for all forecasting scenarios are compiled in Table 1.
Metrics.
@@
ProbTS incorporates a broad range of evaluation metrics to enable a thorough assessment of both point and distributional forecasts. These metrics are elaborated in detail in Appendix A.6. In this paper, we primarily use the normalized mean absolute error (NMAE) for point forecasts and the continuous ranked probability score (CRPS) for distributional forecasts to succinctly communicate the critical insights discovered. It is noteworthy that some methods reproduced in ProbTS, their original papers reported certain point forecast metrics before de-normalizing forecasts to the initial scale (Zeng et al., 2023; Wu et al., 2023; Nie et al., 2023) or primarily reveal aggregated distributional metrics over all time-series variates, namely CRPS-sum (Salinas et al., 2019; Rasul et al., 2021a, b). We have diligently ensured our reproduced results align with their reported results and utilized the unified metrics in this study to offer a comprehensive and fair comparison.
Model.
The design of our model module is modularized to accommodate various methodological focuses, such as point versus distributional forecasts, customized versus general neural architectures for encoding time-series representations, and autoregressive versus non-autoregressive decoding schemes. This design enables us to easily integrate with existing methods by selecting different routes and configuring specific components appropriately. Specifically, our design is grounded on the following formulation for time-series forecasting that considers different forecasting paradigms and decoding schemes.
Formulation.
@@
We denote an element of a multivariate time series as $x^{k}_{t}\in\mathbb{R}$, where $k$ represents the variate index and $t$ denotes the time index. At time step $t$, we have a multivariate vector $\bm{x}_{t}\in\mathbb{R}^{K}$. Each $x^{k}_{t}$ is associated with covariates $\bm{c}_{t}^{k}\in\mathbb{R}^{N}$, which encapsulates auxiliary information about the observations. Given a length-$T$ forecast horizon a length-$L$ observation history $\bm{x}_{t-L:t}$ and corresponding covariates $\bm{c}_{t-L:t}$, the objective in time series forecasting is to generate the vector of future values $\bm{x}_{t+1:t+T}$. In ProbTS, we decouple a model into an encoder $f_{\phi}$ and a forecaster $p_{\theta}$. An encoder is tasked with generating expressive hidden states $\bm{h}\in\mathbb{R}^{D}$. Under autoregressive decoding scheme, encoder forecasts variates using their past values, which can be formulated as $\bm{h}_{t}=f_{\phi}(\bm{x}_{t-1},\bm{c}_{t},\bm{h}_{t-1})$. Under the non-autoregressive scheme, the encoder generates all the forecasts in one step, which can be expressed as $\bm{h}_{t+1:t+T}=f_{\phi}(\bm{x}_{t-L:t},\bm{c}_{t-L:t+T})$. A forecaster $p_{\theta}$ is employed either to directly estimate point forecasts as $\hat{\bm{x}}_{t}=p_{\theta}(\bm{h}_{t})$, or to perform sampling based on the estimated probabilistic distributions as $\hat{\bm{x}}_{t}\sim p_{\theta}(\bm{x}_{t}|\bm{h}_{t})$.
@@
Table 1: This table presents a quantitative assessment of the inherent characteristics for all forecasting scenarios, each corresponding to a dataset with a specific forecasting horizon. We use the suffixes ”-S” and ”-L” to differentiate between short-term and long-term scenarios. Quantified indicators encompass trend and seasonality strengths, as well as non-Gaussianity, where a higher value signifies a greater deviation from a Gaussian distribution.

### v4

# ProbTS: Benchmarking Point and Distributional Forecasting across Diverse Prediction Horizons

@@ Abstract
Delivering precise point and distributional forecasts across a spectrum of prediction horizons represents a significant and enduring challenge in the application of time-series forecasting within various industries. Prior research on developing deep learning models for time-series forecasting has often concentrated on isolated aspects, such as long-term point forecasting or short-term probabilistic estimations. This narrow focus may result in skewed methodological choices and hinder the adaptability of these models to uncharted scenarios. While there is a rising trend in developing universal forecasting models, a thorough understanding of their advantages and drawbacks, especially regarding essential forecasting needs like point and distributional forecasts across short and long horizons, is still lacking. In this paper, we present ProbTS, a benchmark tool designed as a unified platform to evaluate these fundamental forecasting needs and to conduct a rigorous comparative analysis of numerous cutting-edge studies from recent years. We dissect the distinctive data characteristics arising from disparate forecasting requirements and elucidate how these characteristics can skew methodological preferences in typical research trajectories, which often fail to fully accommodate essential forecasting needs. Building on this, we examine the latest models for universal time-series forecasting and discover that our analyses of methodological strengths and weaknesses are also applicable to these universal models. Finally, we outline the limitations inherent in current research and underscore several avenues for future exploration.  [^1] Project repository: https://github.com/microsoft/ProbTS
@@ 1 Introduction
Time-series forecasting has extensive applications in various industries, including traffic flow forecasting [42], renewable energy forecasting [65], and diverse forecasting demands in retail [9], finance [30], and climate [47]. It is crucial to provide forecasts across different prediction horizons, addressing both short- and long-term planning needs [14, 27, 5, 59]. Moreover, modern decision-making processes typically require not only point forecasts to quantify planning efficiency but also robust distributional estimations to manage uncertainty effectively [25, 31]. The fundamental need to produce accurate point and distributional forecasts across various horizons presents significant challenges to existing forecasting approaches.
Nevertheless, much of the previous research on developing deep learning models for time-series forecasting has often focused on isolated aspects, such as long-term point forecasting or short-term distribution estimations. This narrow focus may result in skewed methodological choices and hinder the adaptability of these models to rarely evaluated scenarios. For example, studies such as [74, 70, 38, 73, 12, 69, 48, 71, 39] have primarily explored neural architecture designs tailored for long-term point forecasting with strong trending and seasonal patterns. However, it remains unclear how these advancements can be effectively extended to capture complicated distributions and whether these designs maintain their effectiveness in short-term scenarios. Conversely, research such as [56, 55, 60, 8, 33] adapts deep generative models [18, 29] for probabilistic forecasting, specializing in characterizing complex data distributions. Yet, these models have mainly been developed and evaluated in short-term scenarios, leaving questions about their effectiveness in long-term forecasting and their ability to preserve point forecasting performance.
Despite the recent surge in building time-series foundation models over the past year [19, 54, 15, 11, 16, 26, 40, 21, 68, 3, 24], our understanding of their advantages and limitations, especially regarding essential forecasting needs like point and distributional forecasts across various horizons, is still limited. Many of these models claim to support arbitrary prediction horizons, employing different mechanisms that come with their own set of advantages and drawbacks. Among them, a select few offer capabilities for distributional forecasting, which, however, are typically confined to predefined closed-form distributions [54, 68] or discrete distributions with value quantization [3]. The emergence of these foundation models has brought about unprecedented zero-shot forecasting capabilities. Consequently, it is both timely and crucial to delve into an evaluation of their strengths and weaknesses, especially in relation to the fundamental forecasting needs mentioned earlier.
In this study, we present ProbTS, a benchmark tool crafted to serve as a comprehensive platform for assessing those key forecasting needs and for performing a detailed comparison of several state-of-the-art models developed in recent years. To address the core forecasting requirements, ProbTS includes a broad array of datasets and spans various forecasting horizons. It also utilizes both point and distributional metrics to facilitate a thorough performance evaluation.
Our research reveals that the specific data characteristics inherent to different forecasting requirements often play a crucial role in guiding the selection of model designs. As a result, it is crucial to have a comprehensive view of the essential forecasting needs. To aid in the analysis and interpretation of performance, we measure three essential data characteristics in ProbTS: the strength of trends and seasonality, and the complexity of the data distribution. Moreover, we have explicitly distinguished two fundamental methodological aspects within ProbTS that differentiate the existing forecasting models, largely influencing their pros and cons. The first aspect pertains to the approach for enabling distributional forecasts. This ranges from models focused solely on point forecasting [48, 39], to those employing pre-defined distribution heads based on certain assumptions about the data distribution [54, 68], and to sophisticated neural probabilistic modules designed to flexibly adapt to various data distributions [55, 60]. The second aspect is the decoding scheme used to generate multi-step forecasts, which can be either autoregressive (AR) or non-autoregressive (NAR).
By utilizing ProbTS, we conduct a systematic comparison between studies that focus on long-term point forecasting and those aimed at short-term distributional estimation, employing various forecasting horizons and evaluation metrics. Our overarching finding is that the strengths of these methods tend to diminish in scenarios they are rarely evaluated in, highlighting several important but unresolved research questions. Notably, while recent probabilistic forecasting approaches have shown proficiency in short-term distribution estimation, we discover that long-term distributional forecasting remains a significant challenge. This challenge stems from the difficulty in achieving adaptive distribution estimation that remains both efficient and effective as the prediction horizon extends—a topic that has not been thoroughly investigated in existing literature. Additionally, our analysis uncovers a clear divide in the choice of decoding schemes: most long-term point forecasting methods opt for NAR, whereas choices in short-term forecasting studies are more evenly split. Further investigation suggests that this preference may be linked to the challenge of error accumulation in AR decoding, particularly as the forecasting horizon lengthens or when dealing with strong trends. However, we also discover that AR decoding tends to perform better in scenarios with pronounced seasonality, indicating a potential area for further exploration in decoding strategies, especially for long-term forecasting. Together with the notable inefficiency of existing NAR-based probabilistic methods, such as CSDI [60], for long-term forecasts, our comparison between NAR and AR decoding encourages a wider exploration of decoding strategies in future research.
Furthermore, we have expanded the analytical framework of ProbTS to include an examination of several very recently developed time-series foundation models, which has allowed us to re-validate some of our earlier findings. Interestingly, there appears to be a relatively even split in their preference for AR and NAR decoding schemes. Our analysis reaffirms the limitation of AR in handling time-series data, as we observe that AR-based foundation models tend to excel at shorter horizons. However, their performance advantages often significantly diminish over longer forecasting periods. This underscores the critical need for future research to focus on addressing the issue of error accumulation in AR-based foundation models. Besides, our exploration reveals that current probabilistic foundation models may face challenges when dealing with complex data distributions. This observation suggests that the integration of more sophisticated distribution estimation techniques could enhance the development of time-series foundation models.
In summary, we have made the following contributions.
  • Introduction of ProbTS, a benchmark tool designed for a thorough evaluation of essential forecasting needs, towards precise point and probabilistic forecasting across varied horizons.
  • Comprehensive analysis of methodological variations within forecasting models, especially regarding distributional estimation methods and decoding schemes (AR vs. NAR), which illuminates significant yet previously underexplored research challenges.
  • Extension of our analytical framework to include the latest time-series foundation models, providing insights into the implications of their methodological choices and underscoring important directions for future research endeavors.
@@ 2 Related Work
Classical Time-series Forecasting Models
In recent years, classical research in time-series forecasting has bifurcated into two distinct but complementary streams. The first stream has concentrated on refining neural architecture designs for long-term forecasting, primarily employing non-autoregressive decoding schemes to address scenarios with pronounced trend and seasonality. This stream has evolved from enhancing multi-layer perceptrons [52, 73] to developing specialized recurrent or convolutional neural networks [34, 37], and introducing Transformer-based models [64, 48, 39]. Despite achieving advancements in point forecasts, these efforts mainly capture average future changes, with only a few adopting approaches like quantile regression to partially overcome this limitation [67, 36]. On the other hand, the second stream, probabilistic time-series forecasting specializes in capturing the intricate data distribution of future time series. It encompasses a spectrum of techniques, from utilizing predefined likelihood functions [53, 58] and Gaussian copulas [57, 20] to exploring advanced deep generative models [56, 8]. Unlike the first stream, this branch employs both AR [56, 55] and NAR decoding schemes [60, 8, 33], often utilizing standard neural network architectures to represent time series [17, 56, 55, 8, 20], though some studies propose customized designs [60, 35, 6]. Together, these streams highlight the diverse approaches to forecasting, ranging from point predictions focusing on the mean future variations to probabilistic forecasts that capture the full distribution of future values. In Appendix A.1, we summarize a comparison of these models on the coverage of essential forecasting needs and their methodological preferences.
Universal Time-series Foundation Models
Over the past year, the development of time-series foundation models has greatly accelerated, driven by the success of language foundation models [10]. This wave has seen models such as Lag-Llama [54], TimesFM [16], Timer [40], and Chronos [3] adopting the decoder-only Transformer architecture with an AR decoding scheme. Conversely, models like ForecastPFN [19], MOIRAI [68], TTM [21], and UniTS [24] employ the NAR decoding, often using variable-length placeholders to indicate prediction positions for different horizons. Probabilistic forecasting is less common, with MOIRAI and Lag-Llama integrating pre-defined distribution heads (Student-t for Lag-Llama and a mixture for MOIRAI) while Chronos uses quantized bins to accommodate time-series values and adopts Softmax outputs for distribution approximation. The strategic choice between AR and NAR decoding and the method for distributional estimation highlight distinct trade-offs. For an extensive comparison, see Appendix A.2.
@@
We observe a plethora of toolkits that have been developed for time-series forecasting. These range from those primarily designed for point forecasting, such as Prophet [61], sktime [41], tsai [51], and TSlib [69], to others that incorporate probabilistic forecasting, including GluonTS [2], PyTorchTS [56], PyTorchForecasting222github.com/jdb78/pytorch-forecasting, and NeuralForecast333github.com/Nixtla/neuralforecast. In creating ProbTS, we built upon the foundations laid by tools like PyTorchTS, GluonTS, and TSlib. Our unique contribution is a detailed approach that supports both precise point and probabilistic forecasting over various horizons, and examines methodological differences in forecasting models, especially regarding distributional estimation and decoding schemes (AR vs. NAR). Additionally, ProbTS integrates cutting-edge time-series foundation models, making it a comprehensive benchmark tool for tackling current and future challenges in time-series forecasting. A comparison of ProbTS with existing toolkits, focusing on functionalities and features, is provided in Appendix A.3.
@@ 3 The ProbTS Tool
We aggregate publicly accessible datasets used for both short-term and long-term forecasting. Initial data visualization analyses reveal that the data domains and forecasting horizons significantly influence specific data characteristics within a given forecasting horizon. For instance, many long-term forecasting scenarios exhibit clear trend and seasonality patterns within a forecasting window, while numerous short-term forecasting cases display irregular variations within a short sliding window. Consequently, we have developed quantified indicators, such as trend and seasonality strengths, along with non-Gaussianity to indicate the complexity of data distribution within a forecasting window. Detailed information about dataset statistics, visualization analyses, and quantified measures can be found in Appendix B.1.1B.1.2B.1.3, and B.1.4. The quantified measurements for all forecasting scenarios are compiled in Table 1.
Metrics
@@
ProbTS incorporates a broad range of evaluation metrics to enable a thorough assessment of both point and distributional forecasts. These metrics are elaborated in detail in Appendix B.2. In this paper, we primarily use the normalized mean absolute error (NMAE) for point forecasts and the continuous ranked probability score (CRPS) for distributional forecasts to succinctly communicate the critical insights discovered. It is noteworthy that some methods reproduced in ProbTS, their original papers reported certain point forecast metrics before de-normalizing forecasts to the initial scale [72, 69, 48] or primarily reveal aggregated distributional metrics over all time-series variates, namely CRPS-sum [57, 55, 56]. We have verified our reproduced results align with their reported results and utilized the unified metrics in this study to offer a comprehensive and fair comparison of these studies from different research threads.
Model
@@
The model module in ProbTS explicitly differentiates critical methodological decisions, especially the decoding scheme (AR vs NAR) and the distributional estimation approach. Specifically, we employ the following mathematical formulation. We denote an element of a multivariate time series as $x^{k}_{t}\in\mathbb{R}$, where $k$ represents the variate index and $t$ denotes the time index. At time step $t$, we have a multivariate vector $\bm{x}_{t}\in\mathbb{R}^{K}$. Each $x^{k}_{t}$ is associated with covariates $\bm{c}_{t}^{k}\in\mathbb{R}^{N}$, which encapsulates auxiliary information about the observations. Given a length-$T$ forecast horizon, a length-$L$ observation history $\bm{x}_{t-L:t}$ and corresponding covariates $\bm{c}_{t-L:t}$, the objective in time series forecasting is to generate the vector of future values $\bm{x}_{t+1:t+T}$. To represent point and distributional forecasting in a unified way, here we divide a model into an encoder $f_{\phi}$ and a forecaster $p_{\theta}$. An encoder is tasked with generating expressive hidden states $\bm{h}\in\mathbb{R}^{D}$. Under autoregressive decoding scheme, encoder forecasts variates using their past values, which can be formulated as $\bm{h}_{t}=f_{\phi}(\bm{x}_{t-1},\bm{c}_{t},\bm{h}_{t-1})$. Under the non-autoregressive scheme, the encoder generates all the forecasts in one step, which can be expressed as $\bm{h}_{t+1:t+T}=f_{\phi}(\bm{x}_{t-L:t},\bm{c}_{t-L:t+T})$. A forecaster $p_{\theta}$ is employed either to directly estimate point forecasts as $\hat{\bm{x}}_{t}=p_{\theta}(\bm{h}_{t})$, or to perform sampling based on the estimated probabilistic distributions as $\hat{\bm{x}}_{t}\sim p_{\theta}(\bm{x}_{t}|\bm{h}_{t})$.
@@
Table 1: This table includes a quantitative assessment of the inherent characteristics for all forecasting scenarios, each corresponding to a dataset with a specific forecasting horizon. We use the suffixes "-S" and "-L" to differentiate between short-term and long-term scenarios. Quantified indicators encompass trend and seasonality strengths, as well as non-Gaussianity, where a higher value signifies a greater deviation from a Gaussian distribution.